OpenStudy (anonymous):

3x/4 = 7+ x/6 helpp me please!

OpenStudy (smurfy14):

x=12

OpenStudy (anonymous):

multiply whole equation by 12 on both sides to delete numerators.

OpenStudy (anonymous):

9x=84+2x. 7x=84 x=12.

OpenStudy (anonymous):

where did 84 come from?

OpenStudy (anonymous):

divide 7 from both sides.

OpenStudy (anonymous):

srry but i still dont understand. i only got 9x = 7 + 2x

OpenStudy (anonymous):

then drag 2x to the other side to make it, 9x-2x=7

OpenStudy (anonymous):

so would you have 7x = 7 x=1

OpenStudy (anonymous):

er....84.

OpenStudy (anonymous):

my bad.

OpenStudy (anonymous):

9x-2x=84.*****

OpenStudy (anonymous):

error there.

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

sorry i really dont get this im only in 8th grade.

OpenStudy (anonymous):

so 7x=84. then divide 7 to make it, 7x/7=84/7.

OpenStudy (anonymous):

cancel the 7's.

OpenStudy (anonymous):

what do you multiply to get 84?

OpenStudy (anonymous):

12.

OpenStudy (anonymous):

x/4 and x/6 have least common multiple of 12.

OpenStudy (anonymous):

so multiply whole equation by 12/1.

OpenStudy (anonymous):

yea then i get 36x/4 = 7+ 12x/6

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

w8, yes. but multiply the 7 as well.

OpenStudy (anonymous):

so since 7 is on the 12x/6 side you wud multiply it by 12?

OpenStudy (anonymous):

this is what u r doing to the whole equation. 12((3x/4=7+x/6))

OpenStudy (anonymous):

then, distribute the 12 to each number.

OpenStudy (anonymous):

12(3x/4)=12(7)+12(x/6). do u get that^^^?

OpenStudy (anonymous):

yea. so your answer wud be 36x/4 = 84 + 12x/6

OpenStudy (anonymous):

yup. then u cancel out common factors.

OpenStudy (anonymous):

36/4=9 so 9x. 84 remains the same. 12/6=2 so 2x.

OpenStudy (anonymous):

so your left with 9x =84 + 2x

OpenStudy (anonymous):

yep. then subtract 2x from both sides to get 9x-2x=84.

OpenStudy (anonymous):

subtract from both side then you get 7x=84!

OpenStudy (anonymous):

There ya go!

OpenStudy (anonymous):

x=12

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

glad i could help!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

you dont know how much i appreciated the step by step help!

OpenStudy (anonymous):

hey no problem! That's why I'm here!

OpenStudy (anonymous):

=) can you help me with another problem?

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

its more challenging than the other one.

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

hit me.

OpenStudy (anonymous):

5x/2 - 10 = x/4 - 1

OpenStudy (anonymous):

step 1....

OpenStudy (anonymous):

ok, so the least common multiple of this one would be 4, right?

OpenStudy (anonymous):

yes. b/c they have the same common denominator.

OpenStudy (anonymous):

then you multiply 4 to the whole equation. show me the result.

OpenStudy (anonymous):

gotcha.

OpenStudy (anonymous):

20x/4 - 10 = 4x/4 -4

OpenStudy (anonymous):

u forgot 10.

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

multiply, 4(-10). remember, all the numbers in the equation.

OpenStudy (anonymous):

20x/4-40=4x/4-4.

OpenStudy (anonymous):

u with me?

OpenStudy (anonymous):

sorry. so its 20x/4 - 40 = 4x/4 - 4

OpenStudy (anonymous):

yes it is.

OpenStudy (anonymous):

so now, you would simplify this equation.

OpenStudy (anonymous):

5x-40= x-4

OpenStudy (anonymous):

yup. good job.

OpenStudy (anonymous):

now you have to isolate x.

OpenStudy (anonymous):

woohoo.

OpenStudy (anonymous):

k so now. i have 4x-40 = -4

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

sweet.

OpenStudy (anonymous):

now the other side.

OpenStudy (anonymous):

x= 9

OpenStudy (anonymous):

yay! good job!

OpenStudy (anonymous):

woohoooo!!! high five!!!

OpenStudy (anonymous):

yay. we did it!

OpenStudy (anonymous):

no, u did it! lol.

OpenStudy (anonymous):

lol. you helped me alot!

OpenStudy (anonymous):

lol i'm here for a reason.

OpenStudy (anonymous):

i know you are! are you like a math professor or something?

OpenStudy (anonymous):

thank you so much once again. i really appreciate the help!

OpenStudy (anonymous):

np! glad I could help!

OpenStudy (anonymous):

=)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

can i ask you another question. its different than these. this is my last question.

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

5y - 2 (y+2) = 5

OpenStudy (anonymous):

ok, do u know the distributive property?

OpenStudy (anonymous):

yes i do.

OpenStudy (anonymous):

distribute 2 to y+2.

OpenStudy (anonymous):

2 or -2

OpenStudy (anonymous):

good question. -2. do u know why?

OpenStudy (anonymous):

remember that theres a minus sign behind the 2

OpenStudy (anonymous):

yes. i got it.

OpenStudy (anonymous):

so you wud have.......... 5y (-2y - 4 ) =5

OpenStudy (anonymous):

no, it's 5y-(-2y)-4=5.

OpenStudy (anonymous):

Okay. so then you wud do keep fix change so now u have 7y -4 =5 ?

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

no!!!

OpenStudy (anonymous):

error.

OpenStudy (anonymous):

yeah. that didnt make sense when i tried to solve it.